# DWGS: Enhancing Sparse-View Gaussian Splatting with Hybrid-Loss Depth Estimation and Bidirectional Warping

### v1

@@ Abstract
Novel View Synthesis (NVS) from sparse views remains a core challenge in 3D reconstruction, typically suffering from overfitting, geometric distortion, and incomplete scene recovery due to limited multi-view constraints. Although 3D Gaussian Splatting (3DGS) enables real-time, high-fidelity rendering, it suffers from floating artifacts and structural inconsistencies under sparse-input settings. To address these issues, we propose DWGS, a novel unified framework that enhances 3DGS for sparse-view synthesis by integrating robust structural cues, virtual view constraints, and occluded region completion. Our approach introduces three principal contributions: a Hybrid-Loss Depth Estimation module that leverages dense matching priors with reprojection, point propagation, and smoothness constraints to enforce multi-view consistency; a Bidirectional Warping Virtual View Synthesis method generates virtual training views to impose stronger geometric and photometric constraints; and an Occlusion-Aware Reconstruction component that utilizes depth-difference mask and a learning-based inpainting model to recover obscured regions. Extensive experiments on standard benchmarks—LLFF, Blender, and DTU—show that DWGS achieves a new state-of-the-art, achieving up to 21.13 dB PSNR and 0.189 LPIPS, while retaining real-time inference capabilities. Code is available at: https://github.com/eternalland/DWGS.
(a) Rendered image and depth map comparison between 3DGS and DWGS.
@@
Figure 1: (a) 3DGS outputs (left) vs. DWGS outputs (right) for both image and depth map. (b) presents an efficiency-quality scatter plot comparing DWGS with various baseline methods.
@@ I Introduction
To address the challenges of sparse-view NVS, recent studies have proposed various approaches[6, 7, 8, 9, 10, 11] . NeRF-based methods mitigate overfitting through depth regularization, semantic consistency, and pre-trained priors, but they remain constrained by long training times and slow inference. 3DGS-based methods exploit the flexibility of explicit representations, exploring depth-guided optimization, multi-view consistency constraints, and diffusion model priors, significantly improving efficiency and quality. For instance, depth regularization methods[11] provide geometric priors via monocular depth estimation or Multi-View Stereo (MVS) with virtual view generation and co-regularization enhance multi-view consistency. Nevertheless, existing methods still face limitations in geometric consistency, occlusion handling, and detail recovery under sparse views, particularly in complex scenes, where scale inconsistencies in monocular depth priors and low overlap in MVS can lead to inaccurate geometry.
To tackle these challenges, we propose DepthWarpGS (DWGS), a novel 3DGS-based framework designed for high-quality and efficient novel view synthesis from sparse input views. The framework introduces three key innovations: First, a Hybrid-Loss Depth Estimation module utilizes a dense matcher to extract sufficient initialization points, followed by a novel composite loss integrating reprojection error, point propagation, and Total Variation (TV) smoothness constraints to obtain robust depth estimates, significantly enhancing multi-view consistency. The point propagation constraint specifically strengthens geometric consistency by exploiting indirect correspondences across multiple views (e.g., deriving constraints between points $a$ and $c$ via their common matches to point $b$ in overlapping images). Second, a Bidirectional Warping Virtual View Synthesis strategy is employed to generate photometrically and geometrically consistent virtual views, providing additional supervision signals to improve robustness. Third, an Occlusion-Aware Reconstruction component addresses occlusion challenges through depth-difference-based foreground mask and a dedicated inpainting model, effectively reconstructing missing regions by prioritizing background content. Figure 1 presents a comparison between DWGS and other baseline methods.
@@
  • A unified DWGS framework integrating depth estimation, virtual view synthesis, and occlusion processing for sparse-view 3DGS that achieves efficient NVS.
  • A robust Hybrid-Loss Depth Estimation module is introduced, leveraging dense matching priors and enforcing multi-view geometric consistency through a composite loss that incorporates reprojection, point propagation, and Total Variation smoothness constraints.
  • Bidirectional Virtual View Synthesis, a novel method that generates virtual training views using Bidirectional Warping. This approach imposes additional photometric and geometric constraints, significantly improving reconstruction quality.
  • Occlusion-Aware Reconstruction, an innovative method is presented to effectively reconstruct heavily occluded regions. This is achieved through a depth-difference-based foreground mask mechanism and a dedicated inpainting model, which collectively prioritize the rendering of background content.
Extensive experiments on LLFF, IBRNet, Blender, DTU, and Tanks&Temples show that DWGS achieves new state-of-the-art performance across metrics including PSNR (up to 21.13dB) and LPIPS (as low as 0.189), while maintaining real-time rendering speeds. The framework demonstrates remarkable robustness across a spectrum of settings, from forward-facing to 360° scenes.
Figure 2: DWGS Pipeline. Sparse images are processed via dense matching, structure from motion (SfM), and monocular depth estimation to obtain correspondences, poses, and depth maps. These are fused in a Hybrid-Loss module for robust point-wise depth. Point cloud depths are aligned with monocular depth maps using least-squares optimization to recover scale. Aligned depths and images undergo Bidirectional Warping and interpolation to produce warped images. Occlusion-Aware Reconstruction utilizes an inpainting model (Simple-LAMA) to restore occluded background regions. The 3D Gaussian scene is reconstructed by optimizing a combined loss function comprising color loss and depth loss.
@@ II-1 NeRF-Based NVS Methods
3DGS represents scenes with explicit Gaussian primitives, achieving real-time rendering via differentiable rasterization, surpassing NeRF in efficiency and quality. 3DGS excels with dense views, supporting tasks like text-to-3D generation [22, 23] and dynamic scene modeling [24, 25]. However, in sparse views, 3DGS suffers from insufficient initial point clouds and limited multi-view constraints, leading to floating artifacts and geometry degradation.
@@
Recent studies optimize sparse-view 3DGS with various strategies. FSGS [26] enhances geometry via unpooling and monocular depth priors but suffers from long training times, limiting its applicability in time-sensitive scenarios. GaussianObject [10] proposes structure-prior-aided initialization, needing only 4 views, and uses diffusion models to repair occluded regions. DNGaussian [27] optimizes Gaussian positions with global-local depth normalization, addressing scale inconsistencies. CoherentGS [28] enhances Gaussian consistency using an implicit decoder and smoothness loss, filling occluded regions. CoR-GS [29] suppresses Gaussian field disagreements via co-regularization, improving geometric consistency. LM-Gaussian [30] integrates stereo priors [31] and diffusion models[32] for iterative detail refinement. MVPGS [33] leverages MVS [34] and Vision Transformers (ViTs) [35] to excavate multi-view cues, optimizing initialization and appearance. MCGS [36] introduces a sparse matcher and progressive pruning for multi-view consistency. SCGaussian [37]introduces a hybrid Gaussian representation with structure-consistent optimization and matching priors to enhance 3D consistency. Binocular-Guided 3DGS [38] constrains rendered depths with binocular stereo consistency, improving view consistency. SID[39], as an enhanced version of FSGS, incorporating semantic regularization from DINO-ViT[40] features and local depth constraints. While existing methods have made progress in specific areas, fundamental challenges including monocular depth scale inconsistencies, robust occlusion handling, and high-fidelity detail recovery in complex scenes remain largely open.
@@
Our DWGS method presents a unified approach for high-quality sparse-view synthesis, integrating three core technical contributions that advance beyond previous methods. DWGS leverages dense matching to establish reliable correspondences, while maintaining training efficiency without Gaussian inflation as in FSGS[26]. The proposed Hybrid-Loss Depth Estimation incorporates multi-view geometric constraints and point propagation, offering stronger robustness than SCGaussian[37], which relies only on reprojection error. The Bidirectional Warping method produces more complete virtual views compared to the forward warping used in MVPGS[33], avoiding hole-filling and post-inpainting. Furthermore, the Occlusion-Aware Reconstruction module performs consistent occlusion handling within a single training stage, unlike the complex two-stage diffusion-based refinement in LM-Gaussian[30]. The unified framework jointly addresses depth, consistency, and occlusion challenges under extreme sparsity, demonstrating clear improvements in rendering quality and operational efficiency.
@@ III Methods
The innovative methods employed in DWGS are introduced as follows: III-A. Preliminaries outlines the 3DGS fundamentals. III-B. Dense matching for sufficient initialization, ray-based 3DGS Optimization reduces degrees of freedom to prevent overfitting. III-C. Hybrid-Loss Depth Estimation improves geometric consistency via hybrid loss fusion. III-D. Bidirectional Warping generates virtual views to enhance photometric and geometric constraints. III-E. Occlusion-Aware Reconstruction handles occlusions using depth-difference mask and inpainting model. Figure 2 illustrates the pipeline of the proposed DWGS framework.
@@ III-A Preliminary for 3D Gaussian Splatting
DWGS leverages 3DGS [5] to represent 3D scenes with explicit Gaussian primitives, enabling real-time NVS through differentiable rasterization. Each Gaussian primitive is defined by a set of differentiable parameters: position vector $\mu$, rotation quaternion $q$, scaling matrix $s$, opacity $\alpha$, and spherical harmonic coefficients $sh$ for view-dependent color, denoted as $\mathcal{G}=\{G_{i}:\mu_{i},q_{i},s_{i},\alpha_{i},sh_{i}\}_{i=1}^{P}$. The position and shape of each Gaussian follow a Gaussian distribution, formulated as:
@@ III-B Dense Matching and Ray-Based Gaussian Splatting
Vanilla 3DGS relies on sparse 3D points from COLMAP, which become critically insufficient under sparse-view conditions, leading to initialization failure and overfitting. To address this, two core prior enhancements are introduced.
@@
Figure 3: Hybrid-Loss Depth Estimation pipeline, which first estimates initial point depth from densely matched points using reprojection and point propagation constraints, then filters outliers. During scene training, rendered depth is refined under reprojection, point propagation, and TV smoothness constraints. Point propagation constraint are computed from the common points. Nearest-neighbor view and secondary filter reducing redundant computation.
@@ III-C1 Initial Depth Estimation Stage
$\mathcal{C}(p_{i}^{a},p_{k}^{c})=\begin{cases}(p_{i}^{a},p_{j}^{b})\in\mathcal{M}_{i,j},\\[5.69054pt] (p_{j}^{b^{\prime}},p_{k}^{c})\in\mathcal{M}_{j,k},\\[5.69054pt] \|p_{j}^{b}-p_{j}^{b^{\prime}}\|<d_{NN},\end{cases}$ (7)
@@
where $R_{ik}$ denotes the relative pose from view $i$ to view $k$, see appendix algorithm 1 for details.
@@
(a) Matching points between two SCGaussian images.
(b) Matching points between two DWGS images.
(c) Depth map comparison after applying different outlier filtering methods between SCGaussian and DWGS.
Figure 4: (a)(b) are the matching point maps of SCGaussian and DWGS, respectively. SCGaussian uses Ransac algorithm to filter outliers, while DWGS uses the Outlier Filtering Mechanism. Compared to SCGaussian, DWGS is more capable of removing outliers and retaining more valid points. (c) compares the depth maps of SCGaussian (left) and DWGS (right).
@@ III-C2 Scene Training Stage
$L_{TV}=\sum_{u,v}\left(|\nabla_{x}D_{i}(u,v)|+|\nabla_{y}D_{i}(u,v)|\right)$ (14)
@@ III-D1 Virtual View Generation Strategy
In sparse-view scenarios, particularly with as few as 3 images, 3DGS rendering often suffers from overfitting and severe geometric distortions due to insufficient multi-view constraints, especially when input views are spatially clustered. To mitigate this, we propose generating multiple virtual views via Depth-Image-Based Rendering (DIBR) to impose stronger photometric and geometric constraints during 3DGS optimization. DIBR, also known as 3D Image Warping, is a core computer graphics technique that projects reference images into 3D space using depth maps and reprojects 3D points onto a virtual camera’s image plane. Two primary warping strategies exist: (1) Forward Warping, which is computationally simple but often produces holes and disocclusions, leading to less effective synthesis; and (2) Backward Warping, which provides higher-quality results but requires the depth map of the target view. Our approach integrates both strategies through Bidirectional Warping, effectively combining their advantages. A quantitative comparison between Forward and Bidirectional Warping is provided in Table I.

### v2

# HBSplat: Robust Sparse-View Gaussian Reconstruction with Hybrid-Loss Guided Depth and Bidirectional Warping

@@ Abstract
Novel View Synthesis (NVS) from sparse views presents a formidable challenge in 3D reconstruction, where limited multi-view constraints lead to severe overfitting, geometric distortion, and fragmented scenes. While 3D Gaussian Splatting (3DGS) delivers real-time, high-fidelity rendering, its performance drastically deteriorates under sparse inputs, plagued by floating artifacts and structural failures. To address these challenges, we introduce HBSplat, a unified framework that elevates 3DGS by seamlessly integrating robust structural cues, virtual view constraints, and occluded region completion. Our core contributions are threefold: a Hybrid-Loss Depth Estimation module that ensures multi-view consistency by leveraging dense matching priors and integrating reprojection, point propagation, and smoothness constraints; a Bidirectional Warping Virtual View Synthesis method that enforces substantially stronger constraints by creating high-fidelity virtual views through bidirectional depth-image warping and multi-view fusion; and an Occlusion-Aware Reconstruction component that recovers occluded areas using a depth-difference mask and a learning-based inpainting model. Extensive evaluations on LLFF, Blender, and DTU benchmarks validate that HBSplat sets a new state-of-the-art, achieving up to 21.13 dB PSNR and 0.189 LPIPS, while maintaining real-time inference. Code is available at: https://github.com/eternalland/HBSplat.
(a) Under Sparse-View Reconstruction: Rendered images and depth maps comparison between MCGS and HBSplat.
@@
Figure 1: (a) MCGS outputs (left) vs. HBSplat outputs (right) for both image and depth map. (b) presents an efficiency-quality scatter plot comparing HBSplat with various baseline methods.
@@ I Introduction
To address the challenges of sparse-view NVS, recent studies have proposed various approaches[6, 7, 8, 9, 10, 11, 12]. NeRF-based methods mitigate overfitting through depth regularization, semantic consistency, and pre-trained priors, but they remain constrained by long training times and slow inference. 3DGS-based methods exploit the flexibility of explicit representations, exploring depth-guided optimization, multi-view consistency constraints, and diffusion model priors, significantly improving efficiency and quality. For instance, depth regularization methods[11] provide geometric priors via monocular depth estimation or Multi-View Stereo (MVS) with virtual view generation and co-regularization enhance multi-view consistency. Nevertheless, existing methods still face limitations in geometric consistency, occlusion handling, and detail recovery under sparse views, particularly in complex scenes, where scale inconsistencies in monocular depth priors and low overlap in MVS can lead to inaccurate geometry.
To tackle these challenges, we propose HBSplat, a novel 3DGS-based framework designed for high-quality and efficient novel view synthesis from sparse input views. The framework introduces three key innovations: First, a Hybrid-Loss Depth Estimation module utilizes a dense matcher to extract sufficient initialization points, followed by a novel composite loss integrating reprojection, point propagation, and Total Variation (TV) smoothness constraints to obtain robust depth estimates, significantly enhancing multi-view consistency. The point propagation constraint specifically strengthens geometric consistency by exploiting indirect correspondences across multiple views (e.g., deriving constraints between points $a$ and $c$ via their common matches to point $b$ in overlapping images). Second, a Bidirectional Warping Virtual View Synthesis strategy is employed to generate photometrically and geometrically consistent virtual views, providing additional supervision signals to improve robustness. Third, an Occlusion-Aware Reconstruction component addresses occlusion challenges through depth-difference-based foreground mask and a dedicated inpainting model, effectively reconstructing missing regions by prioritizing background content. Figure 1 presents a comparison between HBSplat and other baseline methods.
@@
  • We introduce a comprehensive framework that seamlessly integrates depth estimation, virtual view synthesis, and occlusion-aware processing into 3DGS to achieve high-quality NVS from sparse inputs.
  • We develop a robust Hybrid-Loss Depth Estimation module that leverages dense matching priors through a composite loss function, enforcing multi-view geometric consistency by combining reprojection, point propagation, and Total Variation constraints.
  • We propose a Bidirectional Warping strategy to generate photometrically and geometrically consistent virtual training views, thereby significantly enhancing reconstruction quality.
  • We design an innovative Occlusion-Aware Reconstruction mechanism that addresses severe occlusions through a depth-difference-based foreground mask and a learning-based inpainting model to effectively recover background content in heavily occluded regions.
Extensive experiments on LLFF, IBRNet, Blender, DTU, and Tanks&Temples show that HBSplat achieves new state-of-the-art performance across metrics including PSNR (up to 21.13dB) and LPIPS (as low as 0.189), while maintaining real-time rendering speeds. The framework demonstrates remarkable robustness across a spectrum of settings, from forward-facing to 360° scenes.
Figure 2: HBSplat pipeline. First, sparse input images are processed by dense matching, SfM, and monocular depth estimation to obtain correspondences, camera poses, and depth maps. The Hybrid-Loss module fuses these inputs to produce robust point-wise depths. Subsequent least-squares optimization aligns the point cloud with monocular depths, recovering the metric scale. The Bidirectional Warping module leverages these aligned depths and images to synthesize novel virtual training views through depth-image warping and interpolation. Simultaneously, the Occlusion-Aware Reconstruction component restores missing background content in occluded regions using learning-based inpainting model (Simple-LAMA) guided by local foreground mask. Finally, the framework reconstructs the 3D Gaussian scene by optimizing a joint loss that combines color and depth supervision.
@@ II-1 NeRF-Based NVS Methods
3DGS represents scenes with explicit Gaussian primitives, achieving real-time rendering via differentiable rasterization, surpassing NeRF in efficiency and quality. 3DGS excels with dense views, supporting tasks like text-to-3D generation [23] and dynamic scene modeling [24, 25]. However, in sparse views, 3DGS suffers from insufficient initial point clouds and limited multi-view constraints, leading to floating artifacts and geometry degradation.
@@
Recent studies optimize sparse-view 3DGS with various strategies. FSGS [26] enhances geometry via unpooling and monocular depth priors but suffers from long training times, limiting its applicability in time-sensitive scenarios. GaussianObject [10] proposes structure-prior-aided initialization, needing only 4 views, and uses diffusion models to repair occluded regions. DNGaussian [27] optimizes Gaussian positions with global-local depth normalization, addressing scale inconsistencies. CoherentGS [28] enhances Gaussian consistency using an implicit decoder and smoothness loss, filling occluded regions. CoR-GS [29] suppresses Gaussian field disagreements via co-regularization, improving geometric consistency. LM-Gaussian [30] integrates stereo priors [31] and diffusion models[32] for iterative detail refinement. MVPGS [33] leverages MVS [34] and Vision Transformers (ViTs) [35] to excavate multi-view cues, optimizing initialization and appearance. MCGS [36] introduces a sparse matcher and progressive pruning for multi-view consistency. SCGaussian [37]introduces a hybrid Gaussian representation with structure-consistent optimization and matching priors to enhance 3D consistency. Binocular-Guided 3DGS [38] constrains rendered depths with binocular stereo consistency, improving view consistency. SID[39], as an enhanced version of FSGS, incorporating semantic regularization from DINO-ViT[40] features and local depth constraints. While existing methods have made progress in specific areas, fundamental challenges including monocular depth scale inconsistencies, robust occlusion handling, and high-fidelity detail recovery in complex scenes have yet to be adequately resolved.
@@
Our HBSplat method presents a unified approach for high-quality sparse-view synthesis, integrating three core technical contributions that advance beyond previous methods. The proposed Hybrid-Loss Depth Estimation incorporates multi-view geometric constraints including point propagation, offering stronger robustness than SCGaussian[37], which relies only on reprojection error. The Bidirectional Warping method produces more complete virtual views compared to the forward warping used in MVPGS[33], avoiding hole-filling and post-inpainting. Furthermore, the Occlusion-Aware Reconstruction module performs consistent occlusion handling within a single training stage, unlike the complex two-stage diffusion-based refinement in LM-Gaussian[30]. The unified framework jointly addresses depth, consistency, and occlusion challenges under extreme sparsity, demonstrating clear improvements in rendering quality and operational efficiency.
@@ III Methods
This section details the HBSplat framework, which is structured as follows: III-A. Preliminaries outlines the 3DGS fundamentals. III-B. Dense matching for sufficient initialization, ray-based 3DGS Optimization reduces degrees of freedom to prevent overfitting. III-C. Hybrid-Loss Depth Estimation improves geometric consistency via hybrid loss fusion. III-D. Bidirectional Warping generates virtual views to enhance photometric and geometric constraints. III-E. Occlusion-Aware Reconstruction handles occlusions using depth-difference mask and inpainting model. Figure 2 illustrates the pipeline of the proposed HBSplat framework.
@@ III-A Preliminary for 3D Gaussian Splatting
HBSplat leverages 3DGS [5] to represent 3D scenes with explicit Gaussian primitives, enabling real-time NVS through differentiable rasterization. Each Gaussian primitive is defined by a set of differentiable parameters: position vector $\mu$, rotation quaternion $q$, scaling matrix $s$, opacity $\alpha$, and spherical harmonic coefficients $sh$ for view-dependent color, denoted as $\mathcal{G}=\{G_{i}:\mu_{i},q_{i},s_{i},\alpha_{i},sh_{i}\}_{i=1}^{P}$. The position and shape of each Gaussian follow a Gaussian distribution, formulated as:
@@ III-B Dense Matching and Ray-Based Gaussian Splatting
Vanilla 3DGS relies on sparse 3D points from COLMAP, which become critically insufficient under sparse-view conditions, leading to initialization failure and overfitting. Although some methods employ MVSNet [41] or VGGT [42] to generate dense point clouds for initialization, they still suffer from multi-stage cascading errors or scale inaccuracies. Here, we customize two simple yet effective core prior enhancements.
@@
Figure 3: Hybrid-Loss Depth Estimation pipeline, which first estimates initial point depth from densely matched points using reprojection and point propagation constraints, then filters outliers. During scene training, rendered depth is refined under reprojection, point propagation, and TV smoothness constraints. Point propagation constraint are computed from the common points. Nearest-neighbor view and secondary filtering reducing redundant computation.
@@ III-C1 Initial Depth Estimation Stage
$\mathcal{C}(p_{i}^{a},p_{k}^{c})=\begin{cases}(p_{i}^{a},p_{j}^{b})\in\mathcal{M}_{i,j},\\[2.84526pt] (p_{j}^{b^{\prime}},p_{k}^{c})\in\mathcal{M}_{j,k},\\[2.84526pt] |p_{j}^{b}-p_{j}^{b^{\prime}}|<d_{NN},\end{cases}$ (7)
@@
where $R_{ik}$ denotes the relative pose from view $i$ to view $k$. Algorithm 1 describes the process of constructing the common point set $\mathcal{C}$.
Algorithm 1 Common Point Set $\mathcal{C}$
0: Matching point pairs $\mathcal{M}_{i,j}={(\mathbf{p}_{i}^{a},\mathbf{p}_{j}^{b})}$ between views $i$ and $j$, $\mathcal{M}_{j,k}={(\mathbf{p}_{j}^{b^{\prime}},\mathbf{p}_{k}^{c})}$ between views $j$ and $k$, pixel distance threshold $\epsilon$;
0: Common point set $\mathcal{C}={(\mathbf{p}_{i}^{a},\mathbf{p}_{k}^{c})}$ for indirect matching between views $i$ and $k$;
1:Construct KD-Trees for View $j$:
2:    Build KD-Tree $\mathcal{T}_{j}$ using all $\mathbf{p}_{j}^{b}$ from $\mathcal{M}_{i,j}$
3:    Build KD-Tree $\mathcal{T}^{\prime}_{j}$ using all $\mathbf{p}_{j}^{b^{\prime}}$ from $\mathcal{M}_{j,k}$
4:Search for Mutual Nearest Neighbors:
5:    Initialize empty set $\mathcal{C}$
6:    for each $\mathbf{p}_{j}^{b}\in\mathcal{M}_{i,j}$ do
7:      Find its nearest neighbor $\mathbf{p}_{j}^{b^{\prime}}$ in $\mathcal{T}^{\prime}_{j}$
8:      Find the nearest neighbor $\mathbf{p}_{j}^{b^{\prime\prime}}$ of that $\mathbf{p}_{j}^{b^{\prime}}$ in $\mathcal{T}_{j}$
9:      if $\mathbf{p}_{j}^{b}=\mathbf{p}_{j}^{b^{\prime\prime}}$ and $|\mathbf{p}_{j}^{b}-\mathbf{p}_{j}^{b^{\prime}}|<\epsilon$ then
10:        Extract corresponding $\mathbf{p}_{i}^{a}$ from $\mathcal{M}_{i,j}$ and $\mathbf{p}_{k}^{c}$ from $\mathcal{M}_{j,k}$
11:        Add pair $(\mathbf{p}_{i}^{a},\mathbf{p}_{k}^{c})$ to $\mathcal{C}$
12:      end if
13:    end for
14:return $\mathcal{C}$
@@
(a) Matching pair images (left, middle) and depth map(right) of SCGaussian.
(b) Matching pair images (left, middle) and depth map (right) of HBSplat.
Figure 4: Visual comparison of matching pair images and depth maps. (a) and (b) respectively show the matching pair images (left, middle) and corresponding depth maps (right) for SCGaussian and HBSplat. SCGaussian relies on the RANSAC algorithm for outlier filtering, whereas HBSplat employs Outlier Filtering Mechanism. The comparison demonstrates that HBSplat more effectively removes outliers while preserving a greater number of valid points.
@@ III-C2 Scene Training Stage
$L_{TV}=\sum_{u,v}\left(|\nabla_{x}D_{i}(u,v)|+|\nabla_{y}D_{i}(u,v)|\right),$ (14)
@@ III-D1 Virtual View Generation Strategy
In sparse-view scenarios, particularly with as few as 3 images, 3DGS rendering often suffers from overfitting and severe geometric distortions due to insufficient multi-view constraints, especially when input views are spatially clustered. To mitigate this, we propose generating multiple virtual views via Depth-Image-Based Rendering (DIBR) to impose stronger photometric and geometric constraints during 3DGS optimization. DIBR, also known as 3D Image Warping, is a core computer graphics technique that projects reference images into 3D space using depth maps and reprojects 3D points onto a virtual camera’s image plane. Two primary warping strategies exist: (1) Forward Warping, which is computationally simple but often produces holes and disocclusions, leading to less effective synthesis; and (2) Backward Warping, which provides higher-quality results but requires the depth map of the target view. Our approach integrates both strategies through Bidirectional Warping, effectively combining their advantages. Quantitative comparison between Forward and Bidirectional Warping is provided in Table I.